GEOMETRICAL MEANING OF ONE GENERALIZATION OF THE

DETERMINANT OF A SQUARE MATRIX by R.Sušanj and M.Radić, Rijeka, Croatia

Abstract. This paper deals with some properties of the determinant of a rectangular matrix which is a generalization of the determinant of a square matrix. Certain geometrical interpretations involving polygons in the plane R^{2} and polyhedra in the space R^{3} are given. Mathematics subject classification: 15A15, 51M25. Keywords: Determinant of a rectangular matrix, area of a polygon, volume of a polyhedron.